Sam's score was 5 less than 3 times Suzie's score. The sum of both scores is 215. What is the solution?
Answer by stanbon(75887)

Sam's score was 5 less than 3 times Suzie's score. The sum of both scores is 215. What is the solution?
------
Equations:
a = 3s - 5
a + s = 215
-----
Substitute for "a" and solve for "s"::
3s-5 + s = 215
4s = 220
------
s = 55 (Suzie's age now)
a = 3s-5 = 160 (Sam's age now: WOW)
Cheers,
Stan H.
===============